Parallelogram ABCD.


1) Diagonals AC and BD meet at E.DE is 8 cm and AC is 13cm.

Questions:

  • how long is BD?
  • how long is AE?
  • how did you solve for the lengths of line segment BD and AE?
  • what property did you apply?
1
Expert's answer
2022-02-02T16:08:59-0500
"BD=2DE=2(8)=16(cm)"

"AE=\dfrac{AC}{2}=\dfrac{13}{2}=6.5(cm)"


The length of line segment BD is twice the length of line segment BE.


The length of line segment AE is one half of the length of line segment AC.


We applied the property that diagonals of a parallelogram bisect each other.